Question 11 What approximate interest rate (compounded annually) will ensure that R8000 is worth R12000 in five years time?
Answer #1

Ans (3) 8.45%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
12000 / 8000 = (1 + r%)^5
1.5 = (1 + r%)^5
r = 8.45%